Robbie is four years older than his sister, Michela. seven years ago, Robbie was three times as old as his sister. How old is Michela now?
Answer by TimothyLamb(4379)

---
x = y + 4
x - 7 = 3(y - 7)
---
x = y + 4
x - 7 = 3y - 21
---
put the system of linear equations into standard form
---
x - y = 4
x - 3y = -14
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= robbie = 13
y= michela = 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php